OpenStudy (anonymous):

1,2,3,4,5=n+1 3,6,10,15=....how i find this formula?

OpenStudy (anonymous):

what does 1,2,3,4,5=n+1 mean ?

OpenStudy (anonymous):

N m 1 2 2 3 3 4 4 5 for N =0,m=N+1 so, we get m=2 N=1,m=N+1 so we get m=3

OpenStudy (anonymous):

0.5(3n+n^2+2)

OpenStudy (anonymous):

Modifying triangular numbers: \( \huge \frac{(n+1)(n+2)}2 \)

OpenStudy (anonymous):

3,6,10 ,15.. ok.lets we check... for foolformath: n=0 ====)(0+1)(0+2)/2=1 not correct for kenneth: n=0=====) not correct to